### v2

# Soft Phonon Charge-Density Wave Formation in the Kagome Metal KV3Sb5

Abstract
A range of of unusual emergent behaviors have been reported in the charge-density wave (CDW) state of the AAV3Sb5 ($A=\penalty 10000\$K, Rb, Cs) kagome metals, including a CDW formation process without soft phonons, which points to an unconventional CDW mechanism. Here, we use inelastic x-ray scattering to show that the CDW in KV3Sb5 forms via phonons that soften to zero energy at the CDW ordering vector ($L$-point) around $T_{\rm CDW}=78$ K. The intensity of soft phonons exhibit a remarkable in-plane anisotropy, extending over a much larger momentum range along $L$-AA relative to $L$-$H$, which leads to diffuse scattering common among AAV3Sb5. Using first-principles calculations, we find that the momentum-dependent electron-phonon coupling (EPC) is peaked at $L$ and exhibits the same in-plane anisotropy as the phonon softening. Conversely, the electronic susceptibility is not peaked at $L$ and shows the opposite in-plane anisotropy. Our findings favor momentum-dependent EPC as the driving mechanism of the CDW in KV3Sb5, with a CDW formation process similar to that of transition metal dichalcogenides.
The kagome lattice consists of corner sharing triangles, and hosts Dirac crossings, van Hove singularities, and flat bands in its electronic structure Mielke (1991). Rich emergent physics can be realized in kagome metals near the van Hove filling, whereby electronic orders such as ferromagnetism, antiferromagnetism, charge bond orders, $d$-wave Pomeranchuk instabilities, charge-density waves (CDW), and unconventional superconducting phases, have been theoretically predicted Yu and Li (2012); Kiesel and Thomale (2012); Wang et al. (2013); Kiesel et al. (2013). Such electronic orders may be realized in the kagome metals AAV3Sb5 ($A=\penalty 10000\$K, Rb, Cs), which exhibit both CDWs and superconductivity Ortiz et al. (2019, 2020); Wilson and Ortiz (2024).
The AAV3Sb5 compounds are made up of kagome lattices of V atoms coordinated by Sb octahedra, forming quasi-2D V3Sb5 layers separated by alkaline metal $A$ atoms [Fig. 1(a)] Wilson and Ortiz (2024). CDWs with a $2\times 2$ in-plane modulation appear below $T_{\rm CDW}=$ 78 K, 104 K and 94 K for the K, Rb, and Cs variants, respectively. The CDW state exhibits a range of unconventional properties, including an anomalous-Hall-effect-like behavior in the absence of local moments Yang et al. (2020); Yu et al. (2021b); Kenney et al. (2021), broken time-reversal symmetry Jiang et al. (2021); Feng et al. (2021); Mielke et al. (2022), electronic nematicity Nie et al. (2022); Grandi et al. (2023), and switchable chiral transport Guo et al. (2022). Superconductivity with $T_{\rm c}\approx 0.9$ K for (K,Rb)V3Sb5 and 2.5 K for CsV3Sb5 emerge inside the CDW phases at ambient pressure Ortiz et al. (2020); Yin et al. (2021); Ortiz et al. (2021), and persist beyond the suppression of CDW order under pressure Yu et al. (2021a); Chen et al. (2021); Du et al. (2021); Zhou et al. (2024). Because the AAV3Sb5 compounds share similar crystal structures, electronic structures, and CDW modulations, their CDWs likely have a common origin. Angle-resolved photoemission spectroscopy measurements show van Hove saddle points just below the Fermi level Kang et al. (2022); Hu et al. (2022b); Zhu et al. (2023), providing evidence that the CDW is driven by the nesting between these saddle points [Figs. 1(c) and (d)] Tan et al. (2021); Park et al. (2021); Lin and Nandkishore (2021); Denner et al. (2021). Alternatively, electron-phonon coupling (EPC) may also play an important role in driving the CDW in AAV3Sb5 Luo et al. (2022); Xie et al. (2022); Liu et al. (2022); He et al. (2024); Gutierrez-Amigo et al. (2024); You et al. (2025).
How a CDW forms upon cooling towards $T_{\rm CDW}$ offers important clues for its underlying mechanism. Whereas both nesting- and EPC-driven CDWs exhibit phonons that soften to zero energy at $T_{\rm CDW}$, phonon softening is typically limited to a small momentum region in the former, and is extended in the latter Chan and Heine (1973); Zhu et al. (2015, 2017). On the other hand, CDWs that result from strong electron correlations, such as in the cuprates, may exhibit phonons that do not soften to zero energy Le Tacon et al. (2013); He et al. (2018). In the case of (Cs,Rb)V3Sb5, phonon softening was not observed Li et al. (2021); Subires et al. (2023), although such a behavior is unlikely to result from electronic correlations, as AAV3Sb5 are weakly correlated metals Zhou et al. (2023). Instead, such an unconventional CDW formation has been attributed to a persistent phason gap Miao et al. (2021), an unusually large soft phonon linewidth that obscures soft phonons Gutierrez-Amigo et al. (2024), a Jahn-Teller-like instability Wang et al. (2022), a condensation of preformed CDW Park and Son (2023); Liu et al. (2025), or strong quantum fluctuations between distinct metastable CDW states Chen et al. (2025). Although the CDW transitions in (Cs,Rb)V3Sb5 exhibit little or no thermal hysteresis Li et al. (2021), they are unambiguously first-order Song et al. (2022); Zhang et al. (2024), which may obscure the intrinsic CDW formation process in AAV3Sb5. As KV3Sb5 exhibits a second-order CDW transition Scagnoli et al. (2024), it is particularly well-suited for probing the CDW mechanism in AAV3Sb5 via measurements of its CDW formation.
Figure 1: (a) The crystal structure of KV3Sb5. (b) The Brillouin zone of KV3Sb5. (c) Calculated electronic structure of KV3Sb5, the arrows mark van Hove points near the Fermi level. (d) The Fermi surface of KV3Sb5 for $k_{z}=\frac{1}{2}$. The arrows represent the nesting between van Hove points at $M$ ($L$). (e) Normalized resistivity $\rho(T)/\rho(\rm{300\penalty 10000\ K})$ of KV3Sb5 single crystals. The inset shows $d\rho/dT$, with a clear anomaly at $T_{\rm CDW}=78$ K. (f) Temperature dependence of elastic x-ray scattering in KV3Sb5 at ${\bf Q}=(3.5,0,0.5)$.
In this work, we use inelastic x-ray scattering (IXS) measurements to show that the CDW in KV3Sb5 forms via phonons that soften to zero energy around $T_{\rm CDW}=78$ K. The phonon softening occurs over a large region in momentum space, akin to the EPC-driven CDW in transition metal dichalcogenides Weber et al. (2011). First-principles calculations show that the in-plane anisotropy of phonon softening matches that of the momentum-dependent EPC, while differing significantly from the calculated electronic susceptibility. These results show that the CDW transition in KV3Sb5 is driven by matrix elements of the EPC, and given the overall similarity of the AAV3Sb5 compounds, such a mechanism is likely also applicable to (Cs,Rb)V3Sb5.
Single crystals of KV3Sb5 were synthesized using a self-flux method Yu et al. (2021b). Resistivity measurements show a CDW transition at $T_{\rm CDW}=78$ K and a residual resistivity ratio (RRR) of $\sim 39$ [Fig. 1(e)], consistent with previous reports Ortiz et al. (2019, 2021). IXS measurements were carried out using the HERIX spectrometer at the Advanced Photon Source, Argonne National Laboratory, with 23.72 keV incident X-rays Said et al. (2020); Toellner et al. (2011). The energy resolution is determined to be $\sim 1.73$ meV by measuring a plexiglass standard. Momentum transfers are indexed in reciprocal lattice units (r.l.u.) of the high temperature $P6/mmm$ structure, with CDW peaks at ${\bf q}=(0.5,0.5,0.5)$ [$L$-point in Fig. 1(b)] and equivalent positions. First-principles calculations were performed using the Quantum ESPRESSO package Giannozzi et al. (2017), using the Perdew, Burke, and Ernzerhof parameterization of the generalized gradient approximation Perdew et al. (1996). The lattice parameters were optimized with a $\Gamma$-centered $12\times 12\times 8$ $\mathbf{k}$-mesh, and different Gaussian smearings were used to simulate different electronic temperatures Souliou et al. (2022); Korshunov et al. (2023); Cao et al. (2023). The phonon spectrum was obtained using density-functional perturbation theory (DFPT) with a $4\times 4\times 2$ $\mathbf{q}$-mesh Giustino (2017). The electron-phonon coupling was calculated using the EPW Noffsinger et al. (2010) package, and the bare electron susceptibility was calculated using a symmetrized Wannier Hamiltonian Marrazzo et al. (2024); Zhi et al. (2022). See the Supplemental Material for further experimental and computational details 1.
Elastic scattering measured at ${\bf Q}=(3.5,0,0.5)$ reveals a gradual onset of the CDW peak intensity below $T_{\rm CDW}$, demonstrating the second-order character of the CDW transition in KV3Sb5 [Fig. 1(f)], consistent with previous results Scagnoli et al. (2024). IXS scans at various temperatures for ${\bf Q}=(1.5,1.5,4.5)$ (an $L$-point) are shown in Fig. 2(a), revealing a band of scattering with a maximum intensity at $\sim 13$ meV that does not change significantly with temperature, and a mode that softens from $\sim 6.5$ meV towards the elastic line, upon cooling from 140 K to $T_{\rm CDW}$. The electron-phonon coupling $\lambda$ is related to the soft mode energy $E_{\rm sm}$ and the amplitude mode energy $E_{\rm am}$ via $\lambda=(E_{\rm am}/E_{\rm sm})^{2}$. For KV3Sb5, our measurements show $E_{\rm sm}\approx 6.5$ meV at 140 K, and Ref. Wu et al. (2022) shows $E_{\rm am}\approx 12$ meV, which lead to $\lambda\approx 3.4$. This value is reasonably large, and similar to the value inferred for CsV3Sb5 Liu et al. (2022).
To quantitatively assess the temperature-evolution of phonons in Fig. 2(a), the IXS spectrum at each temperature is presented in Fig. 2(b), which are fitted to the sum of a resolution-limited elastic peak, three resolution-limited phonon modes with energies between $\sim 9$ meV and $\sim 13$ meV, and a damped harmonic oscillator (DHO) that characterizes the low-lying mode, after convolution with the instrumental resolution. The DHO takes the form:
$S(E)=\frac{1}{1-\exp(-\frac{E}{k_{\rm B}T})}\frac{A}{E_{0}}\frac{2\gamma EE_{0}}{\pi[(E^{2}-E_{0}^{2})^{2}+(E\gamma)^{2}]},$
where $A$ is the intensity, $E_{0}$ is the undamped phonon energy, and $\gamma$ is the damping rate (peak width for $\gamma\ll E_{0}$) Lamsal and Montfrooij (2016); Song et al. (2023); Cao et al. (2023). The damping rate $\gamma$ renormalizes the phonon energy from $E_{0}$ to $E_{\rm ph}=\sqrt{E_{0}^{2}-\gamma^{2}/4}$, with $E_{\rm ph}\rightarrow 0$ as $\gamma/(2E_{0})\rightarrow 1$. Whereas $E_{\rm ph}$ of the three modes between $\sim 9$ meV and $\sim 13$ meV are essentially temperature-independent, the low-lying mode exhibits a strong temperature dependence, softening towards zero energy upon cooling to $T_{\rm CDW}$. Fitting a power law behavior $E_{\rm ph}=[(T-T_{0})/T_{0}]^{\delta}$ yields $T_{0}=81(3)$ K and $\delta=0.56(8)$, in good agreement with a mean-field softening of phonons towards zero energy at $T_{\rm CDW}$, a behavior also found in $2H$-NbSe2 Weber et al. (2011) . Also closely resembling $2H$-NbSe2 Weber et al. (2011), we find $\gamma/(2E_{0})$ of KV3Sb5 approaches unity upon cooling towards $T_{\rm CDW}$ [Fig. 2(d)].
Figure 2: (a) Color map of scattering intensities for ${\bf Q}=(1.5,1.5,4.5)$ at various temperatures and energies. (b) IXS spectra at various temperatures (vertically shifted), with the corresponding fits shown as solid lines. The resolution-limited elastic peak and phonon modes are shown as dashed and solid lines, respectively. The soft phonon mode is modeled using a DHO, and is represented by the shaded area. The scans at 80 K, 85 K and 90 K are scaled by the factors indicated in the figure. In fitting the 80 K and 85 K data, some parameters were constrained 1. (c) Temperature dependence of the fitted phonon energies $E_{\rm ph}$. The solid blue line is a power law fit. (d) Damping ratio $\gamma/(2E_{0})$ of the soft mode in KV3Sb5, compared with the results for $2H$-NbSe2 Weber et al. (2011).
To explore the momentum dependence of soft phonons, points along ALA-L $\bf Q$} from (1.25, 1.75, 4.5) to (1.5, 1.5, 4.5)] and HLHH-L-H $\bf Q$} from (1.44, 1.44, 1.5) to (1.56, 1.56, 4.5)] are selected for IXS scans [dots in Fig. 3(a)]. Soft phonons are observed over a large momentum range along ALA-L, with significantly softer phonons at 85 K relative to 110 K [Figs. 3(b) and (c)]. The ALA-L phonons at both temperatures are well described by the model in Fig. 2(b), with the momentum dependence of $E_{\rm ph}$ for the low-lying mode shown in Fig. 3(e), revealing phonon softening over an extended momentum range of $\sim 0.25$ Å-1 along ALA-L ($\sim 0.5$ Å-1 along ALAA-L-A). On the other hand, soft phonons are limited to a small momentum range along HLHH-L-H, and are already undetectable 0.1 Å-1 away from $L$ (within 0.2 Å-1 along HLHH-L-H). This in-plane anisotropy of the soft phonon intensity has two contributions: (1) an anisotropy in the phonon dispersion, and (2) an anisotropy in the phonon structure factor 1.
To directly visualize the in-plane anisotropy of the soft phonon intensity, the inelastic scattering in Figs. 3(b)-(d) are integrated from -7 meV to 7 meV and the resulting $I_{\rm int}$ are compared in Fig. 3(f), after subtracting resolution-limited elastic peaks from the data. The results from ALA-L are mirrored to LAL-A. Since the soft phonon mode dominates inelastic scattering in this window, $I_{\rm int}$ is a proxy for the intensity of soft phonons. Compared to intensity at the $L$-point (${\bf Q}_{\rm CDW}$), $I_{\rm int}$ drops by $\sim 30$% for $|{\bf Q}-{\bf Q}_{\rm CDW}|=0.1\penalty 10000\$Å along LHL-H, whereas it drops by $\sim 3$% for a similar displacement from ${\bf Q}_{\rm CDW}$ along LAL-A. The highly anisotropic intensity of soft phonons will give rise to thermal diffuse scattering with a similar anisotropy, as schematically represented by the shaded ellipses in Fig. 3(a). Diffuse scattering with such an in-plane anisotropy is found in CsV3Sb5 and RbV3Sb5 Subires et al. (2023), which demonstrates that the AAV3Sb5 compounds exhibit a common anisotropic diffuse scattering above $T_{\rm CDW}$, pointing to a common underlying CDW mechanism.
Figure 3: (a) Schematic of the reciprocal space probed in this work, with the dashed area zoomed in on the right. Dots indicate reciprocal space positions where IXS scans were performed. Color maps of IXS scans along ALA-L at (b) 85 K and (c) 110 K. (d) Color map of IXS scans along HLHH-L-H at 85 K. See Ref. 1 for detailed IXS scans and fits corresponding to (b)-(d). (e) Dispersion of phonon energies $E_{\rm ph}$ along ALA-L at the two temperatures. (f) Integral of inelastic scattering intensities for points along ALAA-L-A and HLHH-L-H, obtained from data in (b)-(d) after a resolution-limited elastic peak is subtracted. The data for LAL-A is mirrored from that of ALA-L.
To shed light on the CDW mechanism in KV3Sb5, we compute the electronic structure and phonon dispersion of KV3Sb5 using first-principles calculations. The electronic structure of KV3Sb5 is shown in Fig. 1(c), with several van Hove points close to the Fermi level (indicated by arrows), consistent with previous results Kato et al. (2022); Luo et al. (2022). The calculated phonon dispersions are shown in Figs. 4(a), using different Gaussian smearings to simulate temperatures approaching $T_{\rm CDW}$ from above Souliou et al. (2022); Cao et al. (2023); Korshunov et al. (2023). The calculated longitudinal acoustic (LA) mode exhibits a clear dip at $L$ and an in-plane anisotropy around $L$: at 0.01 Ry, the LA phonon dispersion along LHL-H is much stiffer compared to LAL-A, and upon decreasing temperature to 0.0095 Ry, the softening of the LA phonon is more prominent along LHL-H compared to LAL-A. We note that a similar anisotropy in phonon dispersion around $L$ is also obtained in calculations that simulate finite temperatures by considering anharmonic effects Ptok et al. (2022). For comparison, the LA phonon at $L$ becomes imaginary at a low temperature (0.001 Ry), and the dip at $L$ disappears upon increasing the Gaussian smearing to 0.1 Ry 1.
Figure 4: (a) Calculated phonon dispersions in KV3Sb5, for Gaussian smearings of 0.01 Ry and 0.0095 Ry. Similar results for 0.1 Ry and 0.001 Ry are shown in the Supplemental Materials 1. (b) The bare susceptibility $\chi_{0}(q)$ in the $q_{z}=\frac{1}{2}$ plane. The strong peak at AA is due to self-nesting in a quasi-2D system. (c) $\bf q$}-dependent EPC $\lambda_{\bf q}$ for the soft phonon mode in the $q_{z}=\frac{1}{2}$ plane, computed at 0.01 Ry.
In order to determine the mechanism that accounts for the in-plane anisotropy of phonon softening in the calculated phonon dispersion, the bare electronic susceptibility (with form factor) $\chi_{0}({\bf q})$ and the $\bf q$}-dependent EPC $\lambda_{\bf q}$ for the LA phonon are computed at $0.01$ Ry and shown in Figs. 4(b) and (c). $\chi_{0}({\bf q})$ exhibits hexagonal ridges around the zone boundary, these ridges are more intense at $H$ than at $L$, and are more extended along LHL-H than LAL-A. The fact that $\chi_{0}({\bf q})$ is not maximized at $L$ and has an in-plane anisotropy opposite to the calculated anisotropy of phonon softening in Fig. 4(a), suggests that the CDW in KV3Sb5 is unlikely to originate from Fermi surface nesting. This is consistent with previous calculations of the Lindhard response function in AAV3Sb5 Kaboudvand et al. (2022). On the other hand, the EPC $\lambda_{\bf q}$ peaks at $L$, with the peak being more elongated along ALA-L relative to HLH-L, consistent with ysthe calculated phonon softening in Fig. 4(a). These findings strongly favor $\bf q$}-dependent EPC matrix element effects over nesting as the dominant driver of the CDW in KV3Sb5, and are corroborated by low temperature (0.001 Ry) calculations that use pressure to suppress the CDW 1.
Our experimental and theoretical results show that the CDW in KV3Sb5 forms via phonon softening towards zero energy, with the soft phonons exhibiting a prominent in-plane anisotropy that is consistent with the momentum-dependent EPC. The CDW formation in KV3Sb5 exhibits phonon softening over an extended momentum region [Fig. 3(e)], a mean-field temperature dependence of $E_{\rm ph}$ [Fig. 2], and a phonon damping rate $\gamma/(2E_{0})$ that gradually increases towards unity at $T_{\rm CDW}$ [Fig. 2(d)], similar to the EPC-driven CDW of $2H$-NbSe2 Weber et al. (2011), further supporting an EPC-driven CDW in KV3Sb5. The common diffuse scattering [Fig. 3(a)] then suggests EPC to be the common mechanism of the CDWs in AAV3Sb5, although electronic effects Eiter et al. (2012); Enzner et al. (2025) may be also at play and contribute to the unconventional properties reported for these materials.
Our findings raise the question as to why phonon softening is not observed in (Cs,Rb)V3Sb5 Li et al. (2021); Subires et al. (2023), for which there are several potential contributing factors. As the CDW transitions in (Cs,Rb)V3Sb5 are first-order Song et al. (2022); Zhang et al. (2024), the transition may occur before the corresponding phonons significantly soften, which contrasts with KV3Sb5 that shows a second-order CDW transition Scagnoli et al. (2024). Compared to KV3Sb5 that shows a single CDW phase in its temperature-pressure phase diagram Du et al. (2021); Kautzsch et al. (2023), CsV3Sb5 has CDWs modulated by $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},\frac{1}{4})$ at ambient pressure Hu et al. (2022a); Stahl et al. (2022); Xiao et al. (2023), and a pressure-induced $(0,\frac{3}{8},\frac{1}{2})$ CDW under pressure Zheng et al. (2022); Stier et al. (2024). In RbV3Sb5, there may also be a pressure-induced CDW distinct from the ambient pressure CDW, although its transport signatures are much weaker compared to CsV3Sb5 Wang et al. (2021); Du et al. (2022b). Such a coexistence of distinct CDW instabilities close in energy may frustrate the system and hinder the softening of phonons at a particular wave vector Chen et al. (2025). Nonetheless, as the CDW-related diffuse scattering in (Cs,Rb)V3Sb5 Subires et al. (2023) are similar to that in KV3Sb5, the presence of soft phonons in these compounds cannot be excluded. As the soft phonons may be weak in some Brillouin zones [e.g. soft phonons in KV3Sb5 were not detected at ${\bf Q}=(3.5,0,0.5)$], a comprehensive search of soft phonons across multiple zones in (Cs,Rb)V3Sb5 will be helpful to elucidate whether phonon softening is common to the CDW formation of AAV3Sb5.
Given the role of EPC in driving the CDW of AAV3Sb5, it may also play a role in the superconductivity that emerges from the CDW state Jiang et al. (2023); You et al. (2025). In such a scenario, the pairing symmetry is expected to be nodeless, for which there has been considerable evidence Duan et al. (2021); Zhou et al. (2024); Mine et al. (2024); Kaczmarek et al. (2025), although there are also reports of unconventional pairing Guguchia et al. (2023); Deng et al. (2024); Zhao et al. (2024). The role of the lattice (and thus EPC) in the superconductivity of AAV3Sb5 is also underscored by the tuning of $T_{\rm c}$ via structural crossovers or transitions under hydrostatic pressure Tsirlin et al. (2022); Yu et al. (2022); Du et al. (2022a). Furthermore, our calculations show that EPC reasonably accounts for superconductivity of KV3Sb5, when its CDW is destabilized under pressure 1.
In conclusion, we find that the CDW in KV3Sb5 forms via phonon softening towards zero energy around $T_{\rm CDW}$, with soft phonons observed over an extended momentum range along ALA-L but a narrow range along HLH-L. The phonon softening of KV3Sb5 is reproduced in first-principles calculations, and is shown to arise from a momentum-dependent EPC. These results evidence a conventional EPC mechanism of the CDW formation in KV3Sb5, similar to that in transition metal dichalcogenides, which likely also applies to other members of the AAV3Sb5 family.
The work at Zhejiang University was supported by the National Key R&D Program of China (No. 2022YFA1402200, 2023YFA1406303), the National Natural Science Foundation of China (No. 12350710785, 12274363, 12274364, 12304175, 12494592), and the Fundamental Research Funds for the Central Universities (Grant No. 226-2024-00068). This research used resources of the Advanced Photon Source, a U.S. Department of Energy (DOE) Office of Science user facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357.
Note: After the submission of this work, soft phonons were also reported for CsV3Sb5 McGuinness et al. (2025), pointing to a common soft-phonon CDW formation in the AAV3Sb5 compounds.